POLARIZATION QUANTUM BEAT SPECTROSCOPY OF THE A^{1}A′′ STATE OF HCF AND DCF

H. FAN, I. IONESCU, C. MUKARAKATE, C. TAO, M. DESELNICU AND \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

To investigate the ^{19}F and ^{1}H nuclear hyperfine structure, Zeeman effect, and spin-orbit interactions in the simplest singlet carbene, we recorded polarization quantum beat spectra (QBS) of ^{r}R_{0}(J) lines of the pure bending transitions 2_{0}^{n} and combination bands 1_{0}^{1}2_{0}^{n} and 2_{0}^{n}3_{0}^{1} in the A^{1}A′′-X^{1}A′ system of HCF and DCF. The spectra were measured under jet-cooled conditions using a pulsed discharge source, both at zero-field and under application of a weak magnetic field (< 30 G). Analysis yielded the nuclear spin-rotation constants (C_{aa}) and weak field Lande g_{aa} factors, from which the a hyperfine constants were estimated. The use of polarization QBS as a probe of spin-orbit interactions in this system will be highlighted.